OpenStudy (anonymous):

For the systems of linear equation in question: x/3-y/5=4 and 3x/4+2y/3=6. Determine how many solutions exist, use either elimination or substitution to find the solutions (if any), graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection.

OpenStudy (anonymous):

x=174/34 & y=-450/37 by sub.. method

OpenStudy (anonymous):

I recommend solving both equations for y first.

OpenStudy (anonymous):

well, i actually got the x-intercept and the y-intercept... but its in fraction

OpenStudy (anonymous):

You really don't need an x or y-intercept to solve this

OpenStudy (anonymous):

i got x = 696/67 and y = -180/67

OpenStudy (anonymous):

i don't

OpenStudy (anonymous):

Nope.....I will show you the first few steps

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

Okay, the steps are a bit more than what you may like, but this is truly the proper way to do it....

OpenStudy (akshay_budhkar):

x=174/34 & y=-450/37 i guess this is the answer...

OpenStudy (anonymous):

Akshay, how would I graph it?

OpenStudy (anonymous):

\[\frac {x}{3} - \frac{y}{5} = 4\] Add y/5 to both sides: \[\frac {x}{3} = 4 + \frac {y}{5} \] Subtract 4 from both sides: \[\frac {x}{3} - 4 = \frac{y}{5}\] Multiply both sides by 5: \[5(\frac {x}{3} - 4) = y\] Simplify to get: \[y = \frac{5x}{3} - 20\]

OpenStudy (anonymous):

That seems like a bit much, but if you do the same thing with the other equation you end up with: \[y = 9 - \frac{9x}{8}\]

OpenStudy (anonymous):

Now if you set the equations equal to each other you will have: y = y

OpenStudy (anonymous):

Now substitute the two equations in for y: \[\frac{5x}{3} - 20 = 9 - \frac{9x}{8} \]

OpenStudy (anonymous):

Now you have something a little easier to work with..... You can just solve for x from there...

OpenStudy (anonymous):

ok, i think i get it...

OpenStudy (anonymous):

\[ x = \frac{696}{67} \approx 10.3881\]

OpenStudy (anonymous):

i'm trying to understand how you got y=9-9x/8

OpenStudy (anonymous):

yeah, that's what i got...

OpenStudy (anonymous):

That's what the the other equation becomes when you solve for y. You posted two equations

OpenStudy (anonymous):

yeah, i posted 2 equations... but how would I graph it, cause the website that my professor issued us, it won't allow me to put in the decimal

OpenStudy (anonymous):

As a matter of fact, now that I think of it, I better way to solve for y would have been to multiply by factors to get rid of the denominators

OpenStudy (anonymous):

ok, nvm.., got it

OpenStudy (anonymous):

To graph it....I'll give you a link to a program you can use to graph it

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

That's geogebra....or you can just use wolfram alpha.... http://www.wolframalpha.com/

OpenStudy (anonymous):

Either way...

OpenStudy (anonymous):

I personally prefer geogebra because it has CAS and many other features

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're most welcome

OpenStudy (anonymous):

By the way, when you're graphing this, here are the equations you should enter: y=9-(9x)/(8) y = (5x)/(3) - 20 And then the graph will give the intersection points

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

If you solve for y for both equations, then set them equal to each other, you will always get it correct. The other two didn't get it right because the substitution method they use doesn't always yield a correct result. The way I've shown you produces a correct result if there is indeed a solution.